# Solving Stochastic Orienteering Problems with Chance Constraints Using a GNN Powered Monte Carlo Tree Search

### v1

@@ IV-D Training
With the goal of replacing the rollout phase in MCTS with a trained model that can estimate both the utility, $Q$, and failure probabilities, $F$, without simulation, a necessary preliminary step is generating data to train the MPNN. To this end, we started by solving randomly generated instances of the SOPCC problem using the MCTS algorithm described in Section III-B. These training instances are complete graphs with rewards and $(x,y)$ coordinates uniformly and randomly distributed between zero and one (see Figure 2) with edge lengths equal to the Euclidean distance. For training, we generated graphs with 20, 30 and 40 vertices.
@@
We trained all of our networks on a budget $B=2$, and a failure probability $P_{f}=0.1$. During training, travel costs along edges were sampled from an exponential distribution with mean equal to the deterministic edge length (same noise model used in our preliminary work [2].) We noticed that any network trained on more than 6,000 problem instances solutions seemingly converged. Importantly, for training we only used data from successful SOPCC solutions, i.e., we excluded cases where the MCTS algorithm produced a solution that violated the budget constraint. We will go over the impact of these decisions in Section V.
@@ V-A Experiments
We experimentally assessed how the architecture we designed predicts the $Q$ and $F$ values to efficiently solve instances of SOPCC. We compare our GNN-MCTS based method with an exact method based on mixed integer linear program [22] (referred to as MILP in the following), as well as the method we formerly presented in [2], referred to as MCTS-SOPCC. For completeness we also compare the GNN-MCTS against the method described in section III-B that was used for training (referred to as MCTS – recall that while it is based on MCTS-SOPCC it has some slight implementation differences as formerly discussed). Our implementation is written in Python 3.11.5 and we used [13] to implement the network based on pre-existing GNN layers. The test machine has an Intel Core i7-10700F at 2.90GHz, 64GB of RAM, and an Nvidia 1660 SUPER GPU. For the MILP based solution, we used the commercial solver Gurobi. The goal is assessing a value network’s ability to learn heuristics and generalize while solving SOPCC instances of varying complexity. Compared to existing solvers, the strength of the GNN-MCTS solver is expected to be speed. Moreover we expect GNN-MCTS to scale much better as the size of problem instances increase. For both MCTS and MCTS-SOPCC comparison, the number of rollouts $S=100$, i.e., $Q$ and $F$ are estimated averaging 100 the results of 100 independent executions. This value was taken from [2]. Both MCTS and MCTS-SOPCC expand the tree 350 times (a value also taken from [2]. In all cases, we set $P_{f}=0.1$.
Figure 3: Time comparison between the different algorithms. GNN-MCTS took always less than half a second, while MILP times averaged 300 to 400 seconds from graph size 20 to 40 (bars cut to save space).
We start by assessing the time performance of the different algorithms. We display these results in Figure 3 for graphs with different number of vertices, and we display averages over 100 different instances. The most evident result from these experiments is how much faster our algorithm is than those existing. Timing performance is increased upwards of 1000% in most instances.
Timing aside, when solving SOPCC instances one has to consider both the reward collected and the failure rate. Accordingly, we define performance through two variables, average reward obtained per solution, $R$, and average rate of failure, $F$. Obviously, a method that obtains a lot of reward, but largely exceeds the failure rate $P_{f}$ is not practically valuable. Hence, a valuable solution is one that stays within the bounds of the constraints while collecting high cumulative rewards. For this test and the following tests, we reused the same test graphs supplied by [2] for their benchmarking in addition to larger sized graphs. Since we trained on instances with $P_{f}=0.1$, we will compare other methods with the same value. In the experiments, as noted in Table I, we see that GNN-MCTS is at times more conservative and achieves lower failure rates, $F$, while incurring limited reward loss. However, in the case of larger, unseen graph sizes, we note larger failure rates. We will touch more on the implications of network generalization in Section V-C.
In Table II we display the performance ratios between GNN-MCTS, the MCTS model from which our GNN was trained, and the MILP solver. This illustrates how truly fast and effective our model is compared to existing methods. We stayed within 93% or better of our MCTS model, showing that this architecture loses less than 10% performance in these cases. Even more impressive, in some cases we out performed the original model. While performance against the MILP was 60% or better with respect to reward, we can expect improvement with optimal training data, which is much more difficult and time consuming to generate. Figure 5 compares the solutions produced by three of the compared algorithms for an SOPCC instance with 40 vertices.
$P_{f}$ GNN-MCTS MCTS MCTS-SOPCC MILP
Reward Time (s) $F$ Reward Time (s) $F$ Reward Time (s) $F$ Reward Time (s) $F$
@@
[…]
TABLE I: Overall results from benchmarks used in [2] plus even larger graphs averaged over 100 trials. Top rows are for results against previously seen budgets. signify tests run with a model trained against size 40 graphs. & are for results against unseen budgets using same sized model. ! is further explained in V-C.
@@
Figure 5: Proposed solutions for three algorithms graph40 test described in Table I. The green node is start; red is goal. The blue represents nodes of varying reward, darker being higher.
@@
[…]
TABLE II: Ratios between our solver and baseline solvers for reward and time averaged over 100 trials. See Table III for unseen graph sizes.
It is also interesting to explore how the solver performs as $P_{f}$ varies. While in most cases model performance characterized by $P_{f}$ was as expected, a limitation of our implementation is that the activation function modeling $F$ is a sigmoid. This aspect is further discussed in Section V-C.
Ultimately, we achieved the results of real-time approximation in a stochastic environment. Figure 4 shows the solution produced by GNN-MCTS while solving an instance of SOPCC associated with a sampling problem in precision agriculture. The blue dots represent locations to visit, with the color intensity indicating the reward (darker means higher reward). The robot starts at the green vertex and must end at the red vertex before running out of budget.
@@ V-B Ablation Study
Previous literature showed that including spatial information into the vertices attributes is essential when assessing the ability of GNN based methods to solve problems like TSP [5]. However, since solutions of SOPCC instances are also influenced by rewards and remaining budget as well as start and goal locations as pointed out in Section IV-C in our implementation, we correspondingly extended the set of attributes. Therefore, this subsection, through an ablation study, we assess the impact of the temporal and spatial attribute additions, novel to this paper, of the one-hot encoding of the currently visited vertex and the start/goal vertices.
@@
In Figure 6, we see that when we remove any of these additional attributes, we incur in a decrease in average reward. In the case of start/end vertex removal, we saw nearly a 50% reward loss. However, the most significant difference is failure rate, which significantly violates the $P_{f}$ constraint making this an infeasible solution. The same figure also shows that without the current vertex attribute we do not violate $P_{f}$, but the average total reward is worse than the original model. While this solution is more conservative, having a lower $F$ rate, our original model doesn’t violate $P_{f}$ while collecting – on average – more reward. For node level classification using GNNs, these findings are relatively important in that GNNs possess the capability to understand time and space with respect to a given problem. Instead of embedding that information elsewhere in the algorithm, we can instead task the GNN with understanding explicitly defined spatiotemporal data.
@@ V-C Generalization
As with every learning based approach, the quality of the solution is influenced by the nature of the dataset used for training. In our work, with the explicit goal of assessing the ability to learn and generalize, during training we kept $P_{f}=0.1$ and $B=2$, and used the same underlying graph topology, i.e., complete graphs with vertices uniformly scattered in the unit square. With these premises, it is essential to assess how the proposed architecture can generalize to instances characterized by new parameters. We experimented generalization from two different angles: generalization of graph sizes and unseen budgets. The ability to generalize with respect to these two parameters is not only important for robustness, but also to expedite training, especially when it comes to graph size, as larger instances require more time to generate solutions with the MCTS algorithm used for training. As pointed out in Section IV-D, during training we only considered graphs with 20, 30, and 40 vertices.
For generalizing with respect to size, we saw in Table I, within a certain size range, equally competitive performance. For example, we pushed our model to the high end of what we have seen in SOPCC literature (e.g., up to 70 vertices) and found that the generalization is quite effective in terms of reward performance. However, due to the sigmoid problem previously described and well known in literature, for larger problem instances the model requires a tuned $P_{f}$ value. This means that, as displayed in Table I, if one feeds the model with a desired $P_{f}$ value, for unseen problem instances we get solutions with failure probabilities significantly exceeding the limit. This problem can however be remedied by feeding the model with a $P_{f}$ value smaller than the desired one. Note that this hyperparametrization process is necessary also when using the exact method [22] and is not a limitation unique to the method we present. Table III shows how by adjusting the $P_{f}$ value in unseen problem instances GNN-MCTS manages to successfully solve the problem matching the originally desired failure rate.
$P_{f}$ $R$ $F$ MCTS $R$ Ratio
[…]
TABLE III: When hyperparameterizing $P_{f}$ on unseen graph sizes, we see the model performance improve to near ground truth, presumably due to the sigmoid activation function.
In the case of budget generalization, we can see in Table I that our method was able to maintain similar time and performance ratios, shown in Table II, for unseen budgets. Given that solving a SOP with a lower budget is more difficult, it makes sense that the more budget that is given, the simpler the task becomes for the model. However, we see our algorithm with noticeably lower $F$ without $R$ sacrifice implying – again – that our method is more conservative without much performance impact.
@@ VI Conclusions and Future Work
In this paper, we have presented a new solution to the SOPCC combining MCTS with a message passing graph neural network. This is accomplished by training a neural network that can replace the time consuming rollout phase in MCTS. Key to the idea of solving SOPCC using MCTS introduced in [2], the trained model simultaneously predicts both the anticipated collected reward, $Q$, as well as the predicted failure probability, $F$ – a feature not found in previous works using value networks with reinforcement learning. Another advantage of the methodology we presented is that during the expansion phase, a single forward pass of the network makes predictions for all children at once, instead of considering one child at a time. This allows the algorithm to efficiently explore a larger fraction of the search space.
Our method has been trained using SOPCC solutions computed by a slightly modified version of the algorithm presented in [2]. Extensive simulations compared this new solution with the training method, our former MCTS algorithm presented in [2], and an exact solution based on MILP. Our findings show that the proposed method is orders of magnitude faster while meeting the failure constraint and retaining a significant fraction of the collected reward, though in some cases this requires tuning some hyperparameters. Our ablation study also corroborates that the vertices attributes we selected for the embedding are critical for the success of the method.
There are numerous venues for current and future work. First, it would be interesting to train the model with solutions to SOPCC instances produced by different heuristics or, possibly, an ensemble of heuristics and assess if this improves the overall performance, especially when it comes to generalize to large problem instances. Next, it will also be interesting to explore whether different vertex or edge attributes would improve the predictive capability of the currently used network. Along the same lines, we will also explore whether altering the structure of the final layers of the network (e.g., substituting the sigmoids) would result in better results. Finally, we are currently implementing the planner in ROS 2 and integrating it into our current autonomous navigation stack for precision agriculture and assessing how it can be used to solve in real time complex data collection problems in the field.

### v2

@@ IV-D Training
To replace the rollout phase in MCTS with a trained model capable of estimating both utility ($Q$) and failure probability ($F$) without simulation, a necessary preliminary step is the generation of training data for the MPNN. To this end, we solved randomly generated instances of the SOPCC using the MCTS algorithm described in Section III-B. These instances are defined over complete graphs, where rewards and $(x,y)$ coordinates are independently and uniformly sampled from the interval $[0,1]$ (see Figure 2), and edge lengths correspond to Euclidean distances. For training, we generated problem instances with 20, 30, and 40 vertices.
@@
We trained all of our networks using a budget of $B=2$ and a failure probability of $P_{f}=0.1$. During training, travel costs along edges were sampled from an exponential distribution with a mean equal to the deterministic edge length (the same noise model used in our preliminary work [2]). We observed that any network trained on more than 6,000 problem instance solutions appeared to converge, with the entire model size being 8.4 MB. Importantly, training data included only successful SOPCC solutions, i.e., cases where the MCTS algorithm produced routes that did not violate the budget constraint. We discuss the impact of these decisions in Section V.
@@ V-A Experiments
We experimentally evaluated the performance of our GNN-MCTS architecture in predicting the $Q$ and $F$ values for efficiently solving SOPCC instances. We compare our method against an exact approach based on mixed integer linear programming [22] (referred to as MILP), as well as our previously proposed method in [2], referred to as MCTS-SOPCC. Additionally, for completeness, we include comparisons against the MCTS method described in Section III-B used during training and sharing similarities with MCTS-SOPCC but with minor implementation differences.
Our implementation is in Python 3.11.5, leveraging [13] for GNN layers. Tests were conducted on a machine with an Intel Core i7-10700F (2.90 GHz), 64 GB RAM, and an Nvidia 1660 SUPER GPU. The MILP solutions were obtained using the commercial solver Gurobi.
The main objective is to assess the value network’s capability to learn heuristics and generalize across SOPCC instances of varying complexity. Compared to existing solvers, GNN-MCTS is expected to offer significant speed advantages and better scalability as problem size increases.
For fair comparison, both MCTS and MCTS-SOPCC use 100 rollouts ($S=100$) to estimate $Q$ and $F$ values by averaging results from 100 independent executions, consistent with [2]. The search trees in these methods are expanded 350 times, also following [2]. In all experiments, the failure probability threshold is set to $P_{f}=0.1$.
GNN-MCTS Timing Splits
Select (s) Expand (s) Rollout (s) Backpropagate (s)
@@
MCTS Timing Splits
Select (s) Expand (s) Rollout (s) Backpropagate (s)
@@
TABLE I: Timing breakdown between MCTS with and without GNN value network over 100 trials.
The results in Table I show graphs with different numbers of vertices and their associated MCTS timing breakdowns. The most evident result from these experiments is how linearly the rollout timing scales with size, and even improves, on average, as graph size grows. This improvement is due to the single forward pass used to approximate utility and failure chance for all nodes in the graph at once. Note that we observe the standard deviation of average rollout times for the GNN-MCTS large due to the single forward pass, which is averaged out over time until the next required forward pass, after all children have been evaluated from a single node. As shown in Table II, we see that the reward performance is very close while saving upwards of 1000% in overall timing, with similar results as complexity scales.
Timing aside, when solving SOPCC instances, one must consider both the reward collected and the failure rate. Accordingly, we define performance through two variables: the average reward obtained per solution, $R$, and the average rate of failure, $F$. Obviously, a method that collects high rewards but significantly exceeds the failure threshold $P_{f}$ is not practically valuable. Hence, a viable solution is one that remains within the constraints while maximizing cumulative rewards. For this test, and those that follow, we reused the same test graphs supplied by [2] for benchmarking, in addition to larger graphs. Since we trained on instances with $P_{f}=0.1$, we compare all methods using the same threshold. As shown in Table II, GNN-MCTS is, at times, more conservative and achieves lower failure rates $F$, while incurring only limited reward loss. However, in the case of larger, previously unseen graph sizes, we observe higher failure rates. We will further discuss the implications of network generalization in Section V-C.
In Table III, we display the performance ratios between GNN-MCTS, the MCTS model used to generate its training data, and the MILP solver. This illustrates how remarkably fast and effective our model is compared to existing methods. We stay within 93% or better of our MCTS model, demonstrating that this architecture loses less than 10% performance in these cases. Even more impressively, in some cases we outperform the original model. While performance against the MILP achieves 60% or better in terms of reward, we expect improvements with optimal training data—which is significantly more difficult and time-consuming to generate. Figure 4 compares the solutions produced by three of the evaluated algorithms for an SOPCC instance with 40 vertices.
$P_{f}$ GNN-MCTS MCTS MCTS-SOPCC MILP
Reward Time (s) $F$ Reward Time (s) $F$ Reward Time (s) $F$ Reward Time (s) $F$
$\text{graph20}_{B=2}$ 0.1 2.762 0.226 4% 2.967 5.092 10% 3.49 10.99 11% 3.414 316.16 12%
$\text{graph30}_{B=2}$ 0.1 5.675 0.442 10% 5.517 43.969 12% 6.433 23.117 7% 6.973 9.949 10%
$\text{graph40}_{B=2}$ 0.1 5.523 0.447 13% 5.75 70.712 11% 8.052 43.698 10% 8.843 457.13 10%
$\text{graph20}_{B=3}$& 0.1 4.27 0.329 4% 4.207 7.003 7% 5.481 26.64 13% 5.404 601.078 14%
$\text{graph30}_{B=3}$& 0.1 6.519 0.483 7% 6.804 59.63 14% 7.934 47.848 11% 8.779 317.093 10%
$\text{graph40}_{B=3}$& 0.1 7.106 0.555 1% 7.343 99.019 13% 10.434 77.774 12% 11.548 600.609 14%
$\text{graph50}_{B=2}$ 0.1 6.952 0.533 18%! 6.422 45.935 10% 8.394 26.06 4% 9.019 600.79 13%
$\text{graph60}_{B=2}$ 0.1 8.889 0.756 30%! 7.429 148.93 11% 10.29 31.031 10% 10.602 601.09 10%
$\text{graph70}_{B=2}$ 0.1 8.462 0.778 40%! 8.381 205.451 12% 11.721 31.222 11% 12.35 601.459 10%
TABLE II: Overall results from benchmarks used in [2] plus even larger graphs averaged over 100 trials. Top rows are for results against previously seen budgets. signify tests run with a model trained against size 40 graphs. & are for results against unseen budgets using same sized model. ! is further explained in V-C.
@@
Figure 4: Proposed solutions for three algorithms graph40 test described in Table II. The green node is start; red is goal. The blue represents nodes of varying reward, darker being higher.
@@
$\text{graph20}_{B=2}$ 0.809 1428.57 0.931 22.73
$\text{graph30}_{B=2}$ 0.814 22.73 1.029 100.00
$\text{graph40}_{B=2}$ 0.625 1111.11 0.961 158.73
$\text{graph20}_{B=3}$ 0.79 2000.00 1.015 21.28
$\text{graph30}_{B=3}$ 0.743 666.67 0.958 123.46
$\text{graph40}_{B=3}$ 0.615 1075.27 0.968 178.57
TABLE III: Ratios between our solver and baseline solvers for reward and time averaged over 100 trials. See Table IV for unseen graph sizes.
It is also interesting to explore how the solver performs as $P_{f}$ varies. While in most cases, model performance characterized by $P_{f}$ was as expected, a limitation of our implementation is that the activation function modeling $F$ is a sigmoid, which does not fully capture failure as a function.
Ultimately, we achieved real-time approximation results in a stochastic environment. Figure 3 shows the solution produced by GNN-MCTS while solving an instance of SOPCC associated with a sampling problem in precision agriculture. The blue dots represent locations to visit, with the color intensity indicating the reward (darker means higher reward). The robot starts at the green vertex and must end at the red vertex before running out of budget.
@@ V-B Ablation Study
Previous literature has shown that including spatial information in vertex attributes is essential for evaluating the effectiveness of GNN-based methods on problems like the TSP [5]. However, because SOPCC solutions are also influenced by rewards, remaining budget, and the start and goal locations, as discussed in Section IV-C, we extended the attribute set accordingly in our implementation. This subsection presents an ablation study to assess the impact of our novel temporal and spatial attribute additions, specifically the one-hot encoding of the currently visited vertex and the start/goal vertices.
@@
In Figure 5, we see that removing any of the additional attributes results in a decrease in average reward. In the case of removing the start/end vertex attribute, we observed nearly a 50% loss in reward. However, the most significant impact is on the failure rate, which exceeds the $P_{f}$ constraint and renders the solution infeasible. The same figure also shows that omitting the current vertex attribute does not lead to a violation of $P_{f}$, but the average total reward is worse than that of the original model. While this solution is more conservative, achieving a lower $F$ rate, our original model maintains compliance with $P_{f}$ while collecting, on average, more reward.
For node-level classification using GNNs, these findings are relatively significant: GNNs can capture temporal and spatial relationships relevant to a given problem. Rather than embedding this information elsewhere in the algorithm, we can instead task the GNN with learning explicitly defined spatiotemporal data.
@@ V-C Generalization
As with any learning-based approach, the quality of the solution depends heavily on the nature of the training dataset. In our work, with the explicit goal of assessing the model’s ability to learn and generalize, we trained using a fixed failure probability $P_{f}=0.1$ and budget $B=2$, employing the same underlying graph topology, i.e., complete graphs with vertices uniformly distributed in the unit square. Given these premises, it is essential to evaluate how well the proposed architecture generalizes to instances characterized by new parameters. We investigated generalization from two perspectives: varying graph sizes and previously unseen budgets. The ability to generalize across these parameters is important not only for robustness but also to accelerate training, particularly with respect to graph size, since larger instances require significantly more time to generate solutions using the MCTS algorithm employed during training. As noted in Section IV-D, our training data consisted solely of graphs with 20, 30, and 40 vertices.
For generalization with respect to size, we observed in Table II that within a certain size range, the model delivers equally competitive performance. For example, we pushed our model to the high end of what is seen in the SOPCC literature (e.g., up to 70 vertices) and found that generalization is quite effective in terms of reward performance. However, due to the sigmoid issue described earlier, for larger problem instances the model requires tuning of the $P_{f}$ value. Specifically, as shown in Table II, when feeding the model a desired $P_{f}$ value for unseen problem instances, the solutions can have failure probabilities that significantly exceed the limit. This issue can be remedied by providing the model with a $P_{f}$ value smaller than the desired one. Note that this hyperparameter tuning process is also necessary when using the exact method [22] and is not a limitation unique to our method. Table IV demonstrates how, by adjusting the $P_{f}$ value in unseen problem instances, GNN-MCTS successfully meets the originally desired failure rate.
$P_{f}$ $R$ MCTS $R$ Ratio $F$
$\text{graph50}_{B=2}$ 0.075 6.171 0.961 6%
$\text{graph60}_{B=2}$ 0.06 7.531 1.014 8%
$\text{graph70}_{B=2}$ 0.075 8.007 0.955 7%
TABLE IV: When hyperparameterizing $P_{f}$ on unseen graph sizes, we see the model performance improve to near ground truth, presumably due to the sigmoid activation function.
In the case of budget generalization, as shown in Table II, our method was able to maintain similar time and performance ratios, detailed in Table III, for unseen budgets. Given that solving a SOP with a lower budget is more difficult, it makes sense that providing more budget simplifies the task for the model. However, we observe our algorithm exhibiting noticeably lower failure rates, $F$, without sacrificing reward, $R$, implying once again that our method is more conservative without a significant performance impact.
@@ VI Conclusions and Future Work
In this paper, we have presented a novel solution to the SOPCC by combining MCTS with a message passing graph neural network. This is achieved by training a neural network to replace the computationally expensive rollout phase in MCTS. Building on the key idea introduced in [2], the trained model simultaneously predicts both the expected collected reward, $Q$, and the predicted failure probability, $F$. This capability was not found in previous works employing value networks with reinforcement learning. Another advantage of our methodology is that, during the expansion phase, a single forward pass of the network predicts values for all child nodes simultaneously, rather than evaluating them one by one. This enables the algorithm to efficiently explore a larger portion of the search space.
Our method was trained using SOPCC solutions generated by a slightly modified version of the algorithm presented in [2]. Extensive simulations compared this new approach against the training method, our previous MCTS algorithm [2], and an exact MILP-based solution. The results demonstrate that our proposed method is orders of magnitude faster while satisfying the failure constraint and preserving a significant portion of the collected reward, though some hyperparameter tuning may be required in certain cases. Our ablation study further confirms that the selected vertex attributes for the embeddings are critical to the method’s success.
There are numerous directions for current and future work. First, it would be interesting to train the model using solutions generated by different heuristics or even an ensemble of heuristics to assess whether this improves overall performance, especially in generalizing to larger problem instances. We also plan to explore applying this model, either in centralized or distributed settings, to multi-robot problems involving communication. Additionally, investigating alternative vertex or edge attributes may enhance the predictive capabilities of the network. Similarly, modifications to the network’s final layers, such as improved activation functions, could yield better results. Finally, we are currently implementing the planner in ROS 2 and integrating it into our autonomous navigation stack for precision agriculture, assessing its real-time applicability to complex data collection problems in the field.